Operations Report – December 23rd

  

Crew 218 Operations Report 23-Dec-19
SOL: 01
Name of person filing report: Luz Maria Agudelo Urrego
Non-nominal systems: NA
Generator: run
Hours run: 13
From what time last night: 1900
To what time this morning: 0800
List any additional daytime hours when the generator was run: N/A
Solar— SOC % (Before generator is run at night: 69 %
Diesel Reading – 87.5%
Station Propane Reading – 80%
Ethanol Free Gasoline: N/A
Water (loft tank) (gal): 26
Water Meter (units): 0146950.2
Water (static tank) (gal): 510
Static to Loft Pump used – Yes
Water in Green Hab (gal): 129
Water in Science Dome (gal): 0
Toilet tank emptied: Yes
Deimos rover used: Still in workshop
Hours:
Beginning charge:
Ending charge:
Currently charging:
Sojourner rover used: Assigned to director
Hours:
Beginning charge:
Ending charge:
Currently charging:
Spirit rover used: Nominal
Hours: 119.2
Beginning charge: (Before EVA): 100%
Ending charge: (On return from EVA, before recharging): 96%
Currently charging: Yes
Opportunity rover used: Still in workshop
Hours:
Beginning charge:
Ending charge:
Currently charging:
Curiosity rover used: Nominal
Hours: 124.4
Beginning charge: 100%
Ending charge: (On return from EVA, before charging): 97%
Currently charging: Yes
Notes on rovers: Opportunity and Deimos off-site for maintenance.
ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): No
Reason for use: N/A
Oil Added? No
# Hours the ATVs were used today: 0
Notes on ATVs: N/A
HabCar used and why, where? Used by Outpost crew for mailing and water
CrewCar used and why, where? No
General notes and comments: After refilling loft tank yesterday, the level went down again. Inspection showed that the hose went deep into the tank, thus causing the water to syphon back to the static tank. The problem has been fixed by pulling the hose to the top of the tank and keeping it in place with duct tape. The tank works properly
Summary of internet: Nominal
Summary of suits and radios: Nominal
Summary of Hab operations: Nominal
Summary of GreenHab operations: routine watering, Martian on greenHab has been capture and released on wild habitat
Summary of Science Dome operations: Nominal
Summary of RAM operations: Nominal
Summary of any observatory issues: Nominal
Summary of health and safety issues: Nominal
Questions, concerns and requests to Mission Support: NA

Luz Ma

ENG

Crew 218

Operations Report – December 23rd

  



Crew 218 Operations Report 23-Dec-19

SOL: 01
Name of person filing report: Luz Maria Agudelo Urrego
Non-nominal systems: NA
Generator: run
Hours run: 13
From what time last night: 1900
To what time this morning: 0800
List any additional daytime hours when the generator was run: N/A
Solar— SOC % (Before generator is run at night: 69 %
Diesel Reading – 87.5%
Station Propane Reading – 80%
Ethanol Free Gasoline: N/A
Water (loft tank) (gal): 26
Water Meter (units): 0146950.2
Water (static tank) (gal): 530
Static to Loft Pump used – Yes
Water in Green Hab (gal): 129
Water in Science Dome (gal): 0
Toilet tank emptied: Yes
Deimos rover used: Still in workshop
Hours:
Beginning charge:
Ending charge:
Currently charging:
Sojourner rover used: Assigned to director
Hours:
Beginning charge:
Ending charge:
Currently charging:
Spirit rover used: Nominal
Hours: 119.2
Beginning charge: (Before EVA): 100%
Ending charge: (On return from EVA, before recharging): 96%
Currently charging: Yes
Opportunity rover used: Still in workshop
Hours:
Beginning charge:
Ending charge:
Currently charging:
Curiosity rover used: Nominal
Hours: 124.4
Beginning charge: 100%
Ending charge: (On return from EVA, before charging): 97%
Currently charging: Yes
Notes on rovers: Opportunity and Deimos off-site for maintenance.
ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): No
Reason for use: N/A
Oil Added? No
# Hours the ATVs were used today: 0
Notes on ATVs: Operated for maintenance
HabCar used and why, where? Used by Outpost crew for mailing and water
CrewCar used and why, where? No
General notes and comments: After refilling loft tank yesterday, the level went down again. Inspection showed that the hose went deep into the tank, thus causing the water to syphon back to the static tank. The problem has been fixed by pulling the hose to the top of the tank and keeping it in place with duct tape. The tank works properly
Summary of internet: Nominal
Summary of suits and radios: Nominal
Summary of Hab operations: Nominal
Summary of GreenHab operations: routine watering, Martian on greenHab has been capture and released on wild habitat
Summary of Science Dome operations: Nominal
Summary of RAM operations: Nominal
Summary of any observatory issues: Nominal
Summary of health and safety issues: Nominal
Questions, concerns and requests to Mission Support: NA

Luz Ma
ENG
Crew 218


	

GreenHab Report – December 23rd

  


Crew 218 GreenHab Report 23-DEC-19

Crew GreenHab Officer: Dr. Jonathan R. Buzan

Environmental control: Heating.

Shade cloth (40% and 30%) on.

Average temperature: 21.9°C, 19%

11:40AM

Floor Unit: 16°C

Electronic: 20.7°C

humidity 20%

1:20PM

Floor Unit: 19°C

Electronic: 23.2°C

humidity 18%

4:30PM

Floor Unit: 17°C

Electronic: 23.9°C

humidity 18%

7:20PM

Floor Unit: 15°C

Electronic: 19.6°C

humidity 20%

Max: 26.0°C, 21%

Min: 17.8°C, 17%

Hours of supplemental light: Light system 7:30PM-12:30AM

Daily water usage for crops: 6.4 Gal.

Daily water usage for research and/or other purposes: N/A

Water in Blue Tank – ~129 Gal.

Time(s) of watering for crops: 7:20 PM

SEASONAL PLANTING HISTORY:

Change to crops: None.

OTHER NOTES:

1st sprouts: N/A

—Martian rodent was captured in trap. Has eaten two pieces of dried sweet potato—Yum!

—Gnats infested strawberry plants. Seen small red ‘arachnid’.

—Leaving door closed at 11:40AM.

—Found a Soil Moisture/Ph Sensor in GreenHab.

—Found an additional Temp/Humidity sensor.

—Reset Martian Rodent Trap with 3 pieces of dried sweet potato in anticipation…

Harvest:

Rosemary 0.5g (for bread); Oregano 0.5g, Basil 3g, Thyme 1g (for tomato sauce) Yum!

Support/supplies requested: Request parsley seeds for future missions.

EVA Report – December 23rd

  

EVA #1

Author: Shefali Rana

Start time: 11.20 AM

End time: 12.16 AM

Narrative: Team conducted a training EVA for Marble Ritual. Crew members collected and loaded rock samples at the collection baskets on Cow Dung Road. CMD and JOU went up the ridge and explored the vantage point. Post survey, CMD, HSO, and JOU came down the trail and dislodged some rocks for analysis. The trip back was led by Spirit followed by Curiosity.

Destination: Marble Ritual

Coordinates (use UTM NAD27 CONUS): 12S 518700 4250800

Participants: Cesare (CMD), Shefali (HSO), Ben (JOU)

Road(s) and routes as per MDRS Map: Cow Dung Road

Mode of travel: Spirit and Curiosity

EVA #2

Author: Shefali Rana

Start time: 12.36 PM

End time: 1.20 PM

Narrative: Team conducted an egress/ ingress training EVA for Marble Ritual. Crew members collected and loaded rock samples at the collection baskets on Cow Dung Road. ENG and GEO went on the trail by the ridge and led the way back to Hab with Spirit leading the way.

Destination: Marble Ritual

Coordinates (use UTM NAD27 CONUS): 12S 518700 4250800

Participants: LuzMa (ENG), Pat (GEO)

Road(s) and routes as per MDRS Map: Cow Dung Road

Mode of travel: Spirit and Curiosity


Shefali Rana

HSO

Crew 218

Commander Report – December 23rd

  



In sim, finally!! This morning, after a few crew photos, we entered simulation. This does not mean that we are “playing pretend” or just enjoying a nice hab in the desert. This means that we are now performing research that we know will enhance and support human exploration of Mars. This means that we are testing and challenging ourselves into keeping isolation, following operational procedures, living in close quarters, and thinking like a Martian. As of today, we are all analog astronaut.

The crew was extremely excited today at performing our first EVAs, including all the difficulties of climbing in the rovers and driving around with all the limitations of the space suits. The feeling is amazing, though. I cannot believe I am here again and I feel so comfortable donning my air pack and helmet once again. Marble Ritual makes us follow a tradition, and I am glad to be part of it and to see my crewmates following it too. Walking over this red, white, brown clay soil inflames the imagination and many times I found myself stopping, looking at the desert landscape and imagining to be on some red remote planet.

Simulation is not only EVA, though, and after coming back and preparing mushroom risotto for the crew, I went around the building checking operations and facilities, before deciding to use my spare time with a short nap, while the rest of the crew was enjoying a well-deserved break after tough days of training.

The news tonight is that we will have precipitation that will make our EVAs impractical, but it also means that we will have a white Christmas, which makes me feel closer to our friends and families on Earth.

Cesare

Commander, MDRS 218

EVA Report – December 23rd

  


EVA #1

Author: Shefali Rana

Start time: 11.20 AM

End time: 12.16 AM

Narrative: Team conducted a training EVA for Marble Ritual. Crew members collected and loaded rock samples at the collection baskets on Cow Dung Road. CMD and JOU went up the ridge and explored the vantage point. Post survey, CMD, HSO, and JOU came down the trail and dislodged some rocks for analysis. The trip back was led by Spirit followed by Curiosity.

Destination: Marble Ritual

Coordinates (use UTM NAD27 CONUS): 12S 518700 4250800

Participants: Crew CMD, HSO, JOU

Road(s) and routes as per MDRS Map: Cow Dung Road

Mode of travel: Spirit and Curiosity

EVA #2

Author: Shefali Rana

Start time: 12.36 PM

End time: 1.20 PM

Narrative: Team conducted an egress/ ingress training EVA for Marble Ritual. Crew members collected and loaded rock samples at the collection baskets on Cow Dung Road. ENG and GEO went on the trail by the ridge and led the way back to Hab with Spirit leading the way.

Destination: Marble Ritual

Coordinates (use UTM NAD27 CONUS): 12S 518700 4250800

Participants: Crew ENG, GEO

Road(s) and routes as per MDRS Map: Cow Dung Road

Mode of travel: Spirit and Curiosity

Shefali Rana

HSO

Crew 218

Journalist Report – December 23rd

  

Journalist Report
Ben Durkee, Crew 218 Journalist

Sol 01

Martians like dehydrated sweet potato!

We awoke this morning to find our trap had worked spectacularly and we had captured the

strawberry bandit. I expected to find something resembling the R.O.U.S.’s (Rodents of

Unusual Size) from the Princess Bride, but the creature was more closely reminiscent of

an Earth field mouse. We attempted to communicate, but the language barrier was too

high to surmount. Alas, maybe we will have more luck in the future with a better

understanding of these creatures (or more sweet potatoes).

Our EVA’s (Extra-Vehicular Activities) today were a hybrid of training and tradition.

We split into a team of three for the first expedition (myself, Cesare and Shefali) and

then Pat and LuzMa paired up for the second. Our task: take the rovers about half a mile

out to Marble Ritual for each crew member to place a pebble in the customary pebble

baskets, as 200 crews have done before us.

5 minutes feels like an eternity in the airlock. Given our lack of internet

connectivity, we can’t really discuss current events. And rock-paper-scissors only

gets you so far. Once we had fully depressurized, we emerged from the Hab and retrieved

our rovers. Shefali and Cesare led the way in Curiosity, and I brought up the rear with

Spirit.

After completing the ceremonial goal of our EVA, we began exploring the surrounding

terrain to build familiarity with the EVA suits. We climbed 60 feet up a nearby hill,

and the view upon cresting it was breathtaking. I only mean that figuratively, thank

goodness. The suits may be a bulky, back-breaking burden, but they do a great job of

maintaining our life support (so I have no place to complain).

I led the charge back to the Hab, after which began the 5 minutes of fun. Once we had

re-pressurized and shed our airtight skin, Pat and LuzMa suited up and set off to follow

in our footsteps. I monitored their journey from the Hab via radio, and we all found

consolation in listening to them struggle with the same things we struggled with.

The dynamic duo returned and we all remained confined to the Hab for the remainder of

the day. After separating to do our individual tasks, some daylight still endured. While

most of the crew relaxed and napped, Pat and I set up Super Smash Brothers for some

future team bonding. Nothing provides more catharsis than digitally beating the stuffing

out of each other.

There seems to be an inclement Martian front on the way that will prohibit us from EVA,

so we are very glad we got to stretch our legs today. We may be exclusively breathing

Hab air for the next few days…

Sol Summary – December 23rd

  

Sol: 01

Summary Title: Waking up on Mars

Author's name: Pat Pesa

Mission Status: Crew is Excited!

Sol Activity Summary: Crew started off the day by taking some pictures before full simulation begins. Once "shutting the hatch" and starting full sim, we went on two EVAs to give the crew a chance to train procedures.

Look ahead plan: Tomorrow we will start weather and structural research projects with an EVA by the CMD, GEO, and ENG. *Note: weather might impact EVA schedule.*

Anomalies in work: None

Weather: Mostly Cloudy, around 30 degrees

Crew physical status: healthy

EVA: Two EVAs both to Marbleridge Wall. The 1st included CMD,HSO,JOU. The 2nd included GEO,ENG.

Reports to be filed: sol summary, commander report, operations report, greenhab report, journalist report, EVA report

Support Requested: none

Pat

Geologist, MDRS 218